# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v1

@@ V Evaluation
As the empirical counterpart of our theoretical investigation, we perform an extensive evaluation by defining a range of scenarios that reflect all possible and realistic combinations between adversarial attacks and defenses. Concretely, the research questions corresponding to the theoretical insights and we would like to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent? 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and can they evade active detection? 3) If yes, can active defenses recoup their performance by also turning adaptive?
Metrics. The first metric we use is the $\ell_{2}$ norm between perturbed and original samples, over ASR (Attack Success Rate), for several reasons: 1) it is significantly more fine grained than a binary indicator of success, thus a better choice for comparing baseline attacks and defenses, their adaptive versions, and to each other; 2) it is not based on an arbitrary perceptual threshold, thus it is more informative of the full range of behavior between perfect attack and perfect defense, 3) a threshold moved around can yield widely varying results, perturbation size is an unaffected global metric, and 4) our RL agents learn optimal policies (also) by using this distance as a signal, where the adversary is rewarded by minimizing it and the defender by maximizing it. The budgets we evaluate over are 1K, 2K and 5K queries. As robustness and classification accuracy are in a natural trade-off, the second metric of interest is the accuracy on benign samples (Clean Acc.) that the original model and the active defense achieve together.
@@ V-A Evaluation Setup
The explicit goal for the agents is – by querying the model and being queried – to learn offensive or defensive policies that are general: they transfer to any other evasion task. Thus after training and validating the agents, the final performance is reported on a fixed hold-out set of 100 adversarial episodes where the starting and original samples are selected at random. As is best practice in AML, candidate samples are only those that are correctly classified by the model. For each scenario we perform a limited hyperparameter and reward function exploration (max 30 trials), with the intention to root out poor combinations rather than exhaust the search space. These are described in more detail in Appendix C.
@@
The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and show the highest evasion potential [32]. BAGS is a search-based methodology and operates in a stochastic manner where every query submitted is a new and potentially better adversarial example. Instead, HSJA is deterministic, composed of 3 different stages, and the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
@@
In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class). Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn. In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content. For the majority of experiments we fix $p=0.5$; we also evaluate adaptive defenses when no attack is present ($p=0$) in subsection V-D.
@@
The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – MNIST and CIFAR10 – and over two models with the same architecture but with different training regimes: one with adversarial training and one without. The transition from single-agent to multi-agent RL hides challenges however. In every scenario, we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [43] agents to discover optimal policies that will also constitute best responses for the full AMG [54]. However, learning independently of other agency breaks the theoretical guarantees of convergence [52] – this can be observed in the sole scenario where both agents learn simultaneously and we thus report them separately: 7 & 8 AA-AD. The full list of scenarios is:
@@
TABLE I: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, for normally and adversarially trained models. Clean Acc. records the accuracy of the base model and any defenses present on benign queries; in the first two scenarios (no active defense) the baseline clean accuracy is recorded. Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks. The green-to-red heatmap is a renormalised range of 0 to 100, where 0 is perfect defense (the gap value) and 100 is perfect attack (lowest value for 5K queries). The asterisk denotes scenarios where input transformations were used for evasion.
@@
    AA-ND – Adaptive Attack / No Defense: The degree at which each attack methodology can be optimized by being adaptive compared to the baseline.
@@
    VA-VD – Vanilla Attack / Vanilla Defense: The performance of our active defense, but the non-adaptive version operating with an empirically defined detection threshold.
@@
    AA-AD – Adaptive Attack / Adaptive Defense: Same scenario as before, but it reports the best-case for the defense instead.
For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [34, 49]. Fixing other policies when computing a best response is representative of learning in cybersecurity environments that can contain a range of agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
@@ V-B Results
For consistency and comparability between evaluations, the listed results are from the same 100 test episodes. The gap value denotes the $\ell_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes. A perfect defense would therefore acquire a value very close to the gap, while a perfect attack would never be able to reduce the perturbation to 00; it can be evaluated, however, relative to the baseline or other attacks. We tested the trained agents on budgets higher than 5K and we discovered that the depicted trends hold; however, to make the agent training tractable and the evaluation wider, we limit the maximum query budget per adversarial episode to 5K. Tables I & II report the results for CIFAR10 and MNIST respectively. The color gradient ranges from the gap value to the minimum value that can be achieved @ 5K queries. Overall, the empirical results help us extract and highlight several important insights, practical observations, general implications, and crucial takeaways for the broader AML field:
TABLE II: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Table I.
@@
1K 2K 5K 1K 2K 5K BAGS HSJA
0: VA-ND 5.28 5.26 5.24 3.63 3.06 2.61 99.37 99.37
1: AA-ND 2.75 2.57 2.47 3.61 3.04 2.58 99.37 99.37
2: VA-VD 8.74 8.28 7.94 5.74 5.73 5.73 99.35 99.20
3: AA-VD 3.39 3.29 3.19 3.58 3.12 2.74 99.33 99.31
4: VA-AD 10.45 10.45 10.45 9.87 9.87 9.87 99.31 99.30
5: AA-TD 3.54 3.27 3.10 4.19 3.54 2.80 99.32 99.20
6: TA-AD 9.85 9.77 9.76 9.99 9.99 9.99 99.28 99.34
7: AA-AD 4.95 4.92 4.92 7.07 6.97 6.76 99.31 99.36
8: AA-AD 10.62 10.62 10.62 8.78 8.71 8.66 99.32 99.36
0: VA-ND 5.28 5.27 5.25 4.65 4.02 3.40 99.15 99.15
1: AA-ND 3.28 3.07 2.94 4.60 3.97 3.35 99.15 99.15
2: VA-VD 8.67 8.44 8.08 5.71 5.70 5.70 99.14 98.89
3 AA-VD 4.14 4.04 3.81 4.69 4.35 3.87 99.13 99.15
4: VA-AD 10.39 10.39 10.39 10.07 10.07 10.07 99.09 99.12
5: AA-TD 4.01 3.80 3.61 4.94 4.44 3.89 99.11 99.12
6: TA-AD 10.05 9.97 9.95 10.20 10.20 10.20 99.07 99.06
7: AA-AD 5.51 5.50 5.50 5.47 5.16 4.99 99.01 99.13
8: AA-AD 10.62 10.60 10.60 10.12 10.12 10.12 99.10 99.01
  • A pervasive observation when comparing the upper and lower halves of each table is that adversarial training adds a small amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
  • Evasive transformations interfere with the attack operation, and only when adaptively controlled together does the attack reach its full potential, as demonstrated by the difference between BAGS and HSJA in CIFAR scenarios 5-8.
  • Adaptive defenses outperform non-adaptive or model hardening, as even against the strongest attacker and the worst case for the defense – the 7th scenario – the perturbation for 5K queries is considerably higher.
  • The initial performance of an attack can be misleading: at first glance HSJA appears to be the better one but it is often outperformed by adaptive BAGS, especially in CIFAR-10 and against active defenses.
  • Not surprisingly, adaptive versions of the vanilla attacks perform better (Scenarios 1 & 0); especially BAGS, as is not well optimized out-of-the-box.
  • The advantage of adaptive attacks however is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
  • The performance of both attacks deteriorates considerably against active defenses, however they reach their full potential only when they are also adaptive.
  • Different attack fundamentals respond differently to active defenses; the gradient-based HSJA is naturally disadvantaged against similarity detection, especially on higher-dimensional data like CIFAR-10.
  • We observe that between consecutive scenarios, where an agent learns against the best opponent policy as previously discovered, the $l_{2}$ values oscillate: following a fixed policy enables the learning of an optimal counter to it.
  • When devising an adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric properties. What did prove effective, however, was pure computation [^1] This reminds us of Sutton’s Bitter Lesson [47]: the observation that progress is often driven by gains in computation rather than problem-specific expert knowledge.: to learn a state embedding from raw input, based on Contrastive Learning. These state embeddings also transfer exceedingly well to different attacks like BAGS.
  • The first time an active defense withstands adaptive attacks is in the 5th CIFAR-10 scenario, for BAGS and HSJA both. This is also the first scenario we had to employ evasive transformations to evade the active defense; subsequently, scenarios 5-8 in CIFAR-10 are the only with evasive transformations.
  • Evasive transformations combine differently with different attacks. All the offensive scenarios in BAGS recoup their performance almost completely. This is not the case for HSJA as the latter requires the preservation of the correct gradient information, something that had largely unsatisfactory solutions even with full white-box access to the model and to the true gradient.
V-C Empirical vs. Adaptive Evasion
So far our evaluations have focused on optimizing attack performance and evasiveness jointly. We are also interested in highlighting and measuring the potential of adaptively controlling the evasive actions without controlling the attack parameters themselves, against our proposed defensive agents. To compare how effective are adaptive evasive actions to the empirically defined ones, we evaluate them along with the approach of Chen et al. [16], configured as described in their work. we keep everything else exactly the same and the attacks themselves use their default, non-adaptive parameters: we want to isolate the effect of evasive actions alone. Table III compares the results between the empirical (yellow) and adaptive control (red) settings, and as we can observe our approach consistently outperforms the empirical definition of evasive transformations. Additionally, we note that the evasive agents were trained only in the upper half of the table, that is against non adversarially trained models: notably, adaptively controlling evasive actions transfers to other models too.
TABLE III: Performance comparison (in $l_{2}$ perturbation) between empirical input transformations (yellow scenarios) and the adaptive control of the same transformations (red scenarios) on CIFAR-10. Bottom half of the table is against adversarially trained models.
Scenario CIFAR-10 Gap: 20.01
BAGS HSJA
1K 2K 5K 1K 2K 5K
BA-VD 8.66 8.44 8.24 10.55 9.93 9.29
CA-VD 8.18 7.92 7.63 8.14 6.76 5.04
BA-TD 8.96 8.70 8.44 11.62 10.85 10.09
CA-TD 8.62 8.26 7.95 11.06 10.46 9.78
BA-AD 16.78 16.07 14.87 10.58 9.88 9.16
CA-AD 9.72 9.30 8.91 10.22 9.73 9.10
BA-VD 8.47 8.26 8.05 10.21 9.76 9.33
CA-VD 7.53 7.11 6.72 9.87 9.10 7.81
BA-TD 8.40 8.21 7.95 11.58 10.88 10.13
CA-TD 7.44 6.78 6.14 11.62 10.70 10.02
BA-AD 16.97 16.13 15.02 10.67 10.00 9.43
CA-AD 9.34 8.66 8.34 10.08 9.62 9.07
V-D Base Rate of Attacks
In all the evaluations so far we use a fixed probability $P(adv)=0.5$ that an incoming query is adversarial. To assess how the active and adaptive defenses (Scenarios 4 & 6) perform in the complete absence of attacks, we evaluate them with $P(adv)=0$ without any retraining; the results are shown in Table IV. We observe a small reduction in the accuracy on clean samples that can be attributed to the considerably different base rate of adversarial and benign queries. Note however that as the probability of each is an intrinsic property of each environment, if the base rate of attacks changes any defensive agent can be retrained to adjust to it.
TABLE IV: Clean accuracy on CIFAR-10 for scenarios 4 & 6, with standard and zero probability of attack. Yellow denotes normally trained models and red adversarially trained.
P(adv) BAGS4 BAGS6 HSJA4 HSJA6
0.5 91.55 91.38 91.59 91.62
0.0 90.91 90.95 90.48 90.86
0.5 87.61 87.50 87.58 87.68
0.0 87.02 87.11 86.70 86.98

### v2

@@ V Evaluation
For evaluation, we define a range of scenarios intended to reflect all possible and realistic combinations between adversarial attacks and defenses, and their adaptive versions. Concretely, the research questions we want to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent? 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and evade active detection? 3) If yes, can active defenses recoup their performance by also turning adaptive?
[Metrics]. We employ ASR (Attack Success Rate) and $\mathbf{L_{2}}$ norm of the perturbation. For the former we set a fixed threshold of 3 for consistency between experiments, while the latter is a more fine-grained metric well suited for comparing baseline attacks, defenses, and their adaptive versions, as it is not based on an (arbitrary) perceptual threshold that can yield widely varying results when moved. The budgets we evaluate over are 1K, 2K and 5K queries. As robustness and classification accuracy are typically in trade-off, the third metric of interest is the benign sample accuracy (Clean Acc.) that the original model and the active defense achieve together.
@@ V-A Evaluation Setup
Our goal is to learn offensive or defensive policies that are general: they transfer to any other evasion task. Thus after training and validating the agents, the final performance is reported on a fixed hold-out set of 100 adversarial episodes where the starting and original samples are selected at random. As is best practice in AML, candidate samples are only those that are correctly classified by the model. For each scenario we perform a limited hyperparameter and reward function exploration (max 30 trials), with the intention to root out poor combinations rather than exhaust the search space, described in more detail in Appendix D.
@@
The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and have the highest evasion potential [23]. BAGS is a stochastic, search-based method where every query submitted is a new and potentially better adversarial example. Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
@@
In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class). Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn. In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content. All experiments are performed with $p=0.5$; we also evaluate our trained defense when no attack is present ($p=0$) in Appendix C.
@@
The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – CIFAR10 and MNIST – and over two models with the same architecture but with different training regimes: one with adversarial training and one without. The transition from single-agent to multi-agent RL hides challenges: we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [42] agents to discover optimal policies that will also constitute best responses for the full AMG [47]. However, learning independently of other agency breaks the theoretical guarantees of convergence [49], eg. in scenarios 7 & 8 where both agents learn simultaneously. The complete list of scenarios is:
@@
TABLE II: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, against normally and adversarially trained models. Clean Acc. reports the accuracy on benign queries of the base model plus any defenses present; in the first two scenarios (no active defense) the baseline clean accuracy is reported. Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks. The asterisk denotes where input transformations were used for evasion.
@@
    AA-ND – Adaptive Attack / No Defense: How more optimal is the adaptive version of an attack compared to the baseline.
@@
    VA-VD – Vanilla Attack / Vanilla Defense: The performance of our active defense, the non-adaptive version that has an empirically defined detection threshold.
@@
    AA-AD – Adaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [31, 38]. Fixing other policies when computing a best response is representative of learning in environments that can contain multiple agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
Comparison to SOTA. In Scenarios 0-8 we evaluate all possible combinations between our attack and our defense. As a baseline to compare to, we additionally evaluate our approach to the state-of-the-art stateful defenses and adaptive attacks, that is Blacklight [23] and OARS [24] respectively. We implement both Blacklight and OARS in our interactive environments by using their publicly available code and parameters. As our environments do not return a rejection signal and to make a fair comparison, for OARS rejection coincides with a non-adversarial decision. We define 5 further scenarios:
  1. 9.
    VA-BD – Vanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
  2. 10.
    OA-BD – OARS Attack / Blacklight Defense: OARS against Blacklight.
  3. 11.
    AA-BD – Adaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
  4. 12.
    OA-TD – OARS Attack / Trained Defense: OARS against our trained defense from Scenario 6.
  5. 13.
    OA-AD – OARS Attack / Adaptive Defense: Our adaptive defense retrained against OARS.
Our experiments were run on multiple machines, however to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for 700 MFLOPs.
@@ V-B Results
For consistency and comparability between evaluations, all results are from the same 100 test episodes. The gap value denotes the $L_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes. By testing the trained agents on budgets higher than 5K we discovered that the trend in reducing $L_{2}$ holds; to make the agent training tractable and the evaluation wider however, we limit the maximum query budget per adversarial episode to 5K. Tables II & III report the results for CIFAR10, while IV reports MNIST. Overall, the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
TABLE III: ASR and mean $L_{2}$ perturbation for CIFAR-10, comparing our Adaptive Attack (AA) and Adaptive Defense (AD) to Blacklight (BD) and OARS (OA).
Adv. Trained Scenario CIFAR-10 Gap: 20.01
BAGS HSJA Clean Acc.
1K 2K 5K ASR 1K 2K 5K ASR BAGS HSJA
*9: VA-BD 9.55 9.32 9.17 0% 8.41 8.19 7.80 15% 91.71 91.71
10: OA-BD 9.46 9.46 9.46 1% 6.54 5.83 4.67 50% 91.71 91.71
11: AA-BD 2.26 1.39 1.32 98% 4.55 3.08 2.44 78% 91.71 91.71
12: OA-TD 20.01 20.01 20.01 0% 7.07 6.38 5.53 50% 91.61 91.59
13: OA-AD 20.01 20.01 20.01 0% 11.03 11.00 10.95 5% 91.61 91.69
*9: VA-BD 9.75 9.56 9.46 0% 8.67 8.50 8.28 7% 87.76 87.76
10: OA-BD 9.79 9.79 9.79 1% 5.77 4.53 3.26 72% 87.76 87.76
11: AA-BD 5.59 4.04 2.55 79% 5.59 4.04 2.55 79% 87.76 87.76
12: OA-TD 20.01 20.01 20.01 0% 6.44 5.49 4.38 65% 87.66 87.64
13: OA-AD 20.01 20.01 20.01 0% 11.31 11.12 10.97 7% 87.66 87.74
  • When comparing the upper and lower halves of each table, we can observe that adversarial training adds a limited amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive stateful (VD), also when transferred (Scen. 12). In HSJA, it reduces ASR by $\sim$90when trained against OA specifically, while it offers similar protection to BD when transferred from another attack.
  • Even against the strongest attacker and for the worst case for the defense (Scen. 7), our AD achieves an ASR $\leq 23\%$.
  • Our adaptive attack (AA) outperforms OARS (Scen. 11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling. and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
  • Evasive transformations interfere with the attack operation, exemplified by the difference between BAGS and HSJA in Scen. 5; for the attack to reach its full potential, they have to be adaptively controlled together.
  • The initial performance of an attack can be misleading: out-of-the-box HSJA appears to be the better attack, but it is often outperformed by adaptive BAGS, especially in CIFAR and against active defenses.
  • In Scen. 1-8, where agents train against the best opponent policy as previously discovered, ASR oscillates as following a fixed policy enables the learning of an optimal counter to it; and eventually plateaus, indicating an equilibrium.
  • The advantage of AA is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
  • The performance of both attacks deteriorates considerably against active defenses, however the defenses reach their full potential only when also adaptive.
  • The first time active defenses effectively resist adaptive attacks is in Scen. 5 of CIFAR; we employed evasive transformations from then onward.
  • Different attack fundamentals respond differently to defenses; the gradient estimation stage of HSJA is naturally disadvantaged against similarity detection, while the jump and binary stages are advantaged.
  • When devising the adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric functioning. What did prove effective, however, was pure computation [^2] This reminds us of Sutton’s Bitter Lesson [50], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we learn an embedding with Contrastive Learning [51] from raw input to use a state, which transfers exceedingly well to other attacks like BAGS.
TABLE IV: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Tables II and III.
@@
1K 2K 5K ASR 1K 2K 5K ASR BAGS HSJA
*0: VA-ND 5.30 5.28 5.26 3% 3.59 3.07 2.61 73% 99.37 99.37
*1: AA-ND 2.74 2.57 2.47 78% 3.61 3.09 2.60 74% 99.37 99.37
*2: VA-VD 7.44 6.66 5.63 22% 5.82 5.78 5.73 2% 99.34 99.20
*3: AA-VD 3.79 3.66 3.44 29% 3.54 3.09 2.77 61% 99.37 99.31
*4: VA-AD 10.57 10.57 10.57 0% 10.05 10.05 10.05 0% 99.31 99.30
*5: AA-TD 3.57 3.29 3.14 39% 5.00 3.97 3.38 36% 99.32 98.84
*6: TA-AD 10.62 10.62 10.62 0% 10.23 10.23 10.18 0% 99.28 99.34
*7: AA-AD 4.89 4.89 4.86 8% 5.06 4.76 4.38 36% 99.31 99.35
*8: AA-AD 10.62 10.62 10.62 0% 10.21 10.21 10.21 0% 99.32 99.23
*9: VA-BD 10.62 10.62 10.62 0% 5.65 5.65 5.65 2% 99.37 99.37
10: OA-BD 10.62 10.62 10.62 0% 4.53 4.00 3.15 46% 99.37 99.37
11: AA-BD 3.83 3.69 3.60 17% 4.18 3.66 3.19 52% 99.37 99.37
12: OA-TD 10.62 10.62 10.62 0% 10.21 10.20 10.20 0% 99.22 99.28
13: OA-AD 10.62 10.62 10.62 0% 10.21 10.20 10.20 0% 99.32 99.28
*0: VA-ND 5.26 5.25 5.24 2% 4.61 4.04 3.41 30% 99.15 99.15
*1: AA-ND 3.28 3.08 2.96 51% 4.59 3.97 3.35 34% 99.15 99.15
*2: VA-VD 7.70 6.86 5.86 17% 5.81 5.78 5.76 2% 99.14 99.12
*3 AA-VD 4.18 4.08 3.86 22% 4.63 4.27 3.86 25% 99.13 99.15
*4: VA-AD 10.55 10.55 10.55 0% 10.02 10.02 10.02 0% 99.09 99.08
*5: AA-TD 4.04 3.74 3.54 27% 5.82 5.09 4.26 16% 99.11 98.78
*6: TA-AD 10.62 10.62 10.62 0% 10.20 10.20 10.20 0% 99.06 99.06
*7: AA-AD 5.59 5.56 5.56 5% 5.47 5.16 4.99 14% 99.09 99.13
*8: AA-AD 10.62 10.62 10.62 0% 10.12 10.12 10.12 0% 99.10 99.01
*9: VA-BD 10.62 10.62 10.62 0% 5.65 5.64 5.64 1% 99.15 99.15
10: OA-BD 10.62 10.62 10.62 0% 5.18 4.80 4.11 17% 99.15 99.15
11: AA-BD 4.31 4.07 3.96 13% 5.04 4.65 4.20 19% 99.15 99.15
12: OA-TD 10.62 10.62 10.62 0% 10.26 10.26 10.26 0% 99.00 99.06
13: OA-AD 10.62 10.62 10.62 0% 10.26 10.26 10.26 0% 99.10 99.06